# Tokenization Matters! Degrading Large Language Models through Challenging Their Tokenization

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable capabilities in language understanding and generation. Nonetheless, it was also witnessed that LLMs tend to produce inaccurate responses to specific queries. This deficiency can be traced to the tokenization step LLMs must undergo, which is an inevitable limitation inherent to all LLMs. In fact, incorrect tokenization is the critical point that hinders LLMs in understanding the input precisely, thus leading to unsatisfactory output. To demonstrate this flaw of LLMs, we construct an adversarial dataset, named as ADT (Adversarial Dataset for Tokenizer), which draws upon the vocabularies of various open-source LLMs to challenge LLMs’ tokenization. ADT consists of two subsets: the manually constructed ADT-Human and the automatically generated ADT-Auto. Our empirical results reveal that our ADT is highly effective on challenging the tokenization of leading LLMs, including GPT-4o, Llama-3, Qwen2.5-max and so on, thus degrading these LLMs’ capabilities. Moreover, our method of automatic data generation has been proven efficient and robust, which can be applied to any open-source LLMs. To the best of our knowledge, our study is the first to investigating LLMs’ vulnerability in terms of challenging their token segmentation, which will shed light on the subsequent research of improving LLMs’ capabilities through optimizing their tokenization process and algorithms.
@@ 1 Introduction
In the last two years, Large Language Models (LLMs) have demonstrated remarkable capabilities in many tasks of artificial intelligence (AI), including natural language generation hoffmann2022training ; nijkamp2023codegen ; zeng2023glm130b , knowledge utilization chowdhery2022palm ; izacard2022atlas , and complex reasoning wei2023chainofthought ; zhou2023leasttomost ; kojima2023large ; taylor2022galactica . Given these capabilities, LLMs have been effectively employed by various application domains, such as healthcare tang2023does ; li2023chatdoctor ; jeblick2022chatgpt , education susnjak2022chatgpt ; bordt2023chatgpt ; Malinka_2023 , law yu2022legal ; nay2023law and so on.
Nonetheless, LLMs’ disadvantages have also been witnessed, including hallucination openai2023gpt ; bang2023multitask ; lin2022truthfulqa , knowledge recency dai2022knowledge ; Kernbach_2022 , and so on. Particularly, we also observed that for some specific queries, LLMs often produce unsatisfactory responses with words that are nonsensical, as illustrated by the two instances in Figure 1. Through checking the LLM’s tokenization results for the input sentences, we found that it is misaligned with human’s correct comprehension for the sentence. Notably, our empirical studies found that this flaw not only exists in some specific LLMs, but also is a universal issue across many mainstream LLMs. We have evaluated several prominent open-source and closed-source LLMs, including Chatglm3 zeng2023glm130b , Baichuan2 yang2023baichuan , Qwen2.5-max  [^1] https://help.aliyun.com/zh/dashscope/developer-reference/model-introduction?spm=a2c4g.11186623.0.i3, and GPT-4o  [^2] https://platformopenai.com/docs/models/gpt4o. Our experiment results reveal that regardless of LLMs’ scales or their claimed capabilities, they inevitably generate incorrect or entirely nonsensical responses for some specific inputs when their tokenization results for the input sentences are obviously wrong. Consequently, we believe that LLMs’ tokenization errors prevent them from accurately understanding the input text, leading to their incorrect responses.
@@
Figure 1: Two instances of LLM generating incorrect response due to incorrect tokenization. Case 1 is a Chinese input instance, of which the English translation is noted below according to its correct tokenization. In Case 2, a space is omitted between ‘moves’ and ‘table’, causing the LLM’s incorrect tokenization, which is detailed in Section 3.2.
@@
As we know, LLMs’ tokenization flaws stem from the algorithms of their tokenizers, most of which are based on subword-level vocabularies. The popular tokenization algorithms include Byte-Pair Encoding (BPE) sennrich2016neural , WordPiece Schuster_Nakajima_2012 , and Unigram kudo2018subword . However, no vocabulary can perfectly cover all possible ways of various expressions in the inputs. When applying the greedy principle, the algorithms are easy to generate unsatisfactory results which are not fully aligned with the correct intention of users’ input. Unfortunately, in cases of tokenization errors, all subsequent optimization operations for LLMs cannot completely solve this underlying problems caused by their tokenization algorithms.
@@
In the domain of natural language processing (NLP), the existing studies related to tokenization primarily focus on refining or optimizing various tokenization algorithms. Meanwhile, the discussions on LLMs’ vulnerability including attack or challenge techniques, have been more concerned with the security of LLMs. About the challenges posed by tokenization deficiencies, only Sander and Max land2024fishing have discussed this issue from the perspective of under-trained tokens in LLMs. It is worth noting that, to the best of our knowledge, there has been no research specifically focusing on the unsatisfactory token segmentation of LLMs, which is indeed a critical concern causing LLMs’ vulnerability.
@@
In this paper, we focus on the critical flaw in LLMs’ tokenization process, and try to reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses for some specific queries. To this end, for the first time, we construct a dataset, namely ADT (Adversarial Dataset for Tokenizer), to challenge the tokenization of different representative LLMs. Our ADT dataset consists of two subsets: the manually constructed ADT-Human and automatically generated ADT-Auto. At first, we export the vocabularies from multiple mainstream LLMs, based on which ADT-Human is constructed. Our experiment results demonstrate that ADT-Human can effectively challenge the LLMs’ tokenization, leading to their completely incorrect responses. Furthermore, we also develop a framework for automatically generating adversarial data to construct the dataset more efficiently. Initially, we export an LLM’s vocabulary and identify the trap words that can influence the model’s performance. By inputting these trap words into GPT-4 openai2023gpt with prompt, we leverage its capability to get available instances, which are then inspected manually to ensure their quality. With minimal human effort, we successfully construct ADT-Auto with 231 instances, validating the effectiveness of our framework and highlighting the inevitable flaws in LLMs’ tokenization once again.
@@
    For the first time, we investigate LLMs’ vulnerability for some special inputs in terms of challenging their token segmentation, providing a new perspective of studying LLMs’ disadvantages.
@@
    Our experiment results obviously reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses, which can shed light on the subsequent work of improving LLMs through optimizing their tokenization.
@@ 2 Related work
Tokenization is a basic but crucial step in NLP. In order to get the models to better understand the text, great improvements have been made to the tokenization process. Currently, the mainstream tokenization approachs are based on sub-word units, including Byte Pair Encoding (BPE), WordPiece, and Unigram. BPE sennrich2016neural forms an extensive vocabulary starting from the base of character-level tokens, merging these token pairs and intercalating them into the vocabulary. The merging rule is to select the adjacent pairs with the highest word frequency. LLMs including GPT-3 brown2020language , RoBERTa liu2019roberta , BART lewis2019bart , Llama2 touvron2023llama , and Baichuan yang2023baichuan are based on BPE. WorldPiece Schuster_Nakajima_2012 is similar to BPE, but it differs in the strategy of merging pairs with reference to mutual information rather than frequency. LLMs built upon WordPiece include BERT devlin2019bert , DistilBERT sanh2020distilbert , and Electra clark2020electra . Different from these two algorithms, Unigram kudo2018subword starts with a large vocabulary and gradually trims it down to a smaller one. It measures the importance of subwords by computing the loss incurred by deleting different subwords, retaining those with high significance. LLMs utilizing Unigram include AlBERT lan2020albert and mBART liu2020multilingual . Besides, many tokenization tools integrating these algorithms have been developed, such as Google’s SentencePiece kudo2018sentencepiece and OpenAI’s tiktoken [^3] https://github.com/openai/tiktoken, which simplify the process of LLMs tokenization greatly. However, there has been no work concerning the shared risks inherent in these algorithms. So we firstly focus on this issue, discussing the underlying risks of LLMs’ tokenization.
@@
With the growing prominence of LLMs, the security and vulnerability of these models have attracted significant attention, and even advanced LLMs like GPT-4 are no exception. A surge of research in this field is underway, with researchers launching attacks on LLMs from various aspects esmradi2023comprehensive ; chowdhury2024breaking including but not limited to, Prompt Injection choi2022prompt , Model Theft krishna2020thieves , Data Reconstruction carlini2021extracting , Data Poisoningwallace2021concealed ; xu2024instructions , and Member Inference Attack 7958568 . For instance, Prompt Injection Attack refers to a scenario where an attacker crafts malicious prompts with the intention of deceiving a language model to generate outputs inconsistent with its training data and anticipated functionality. Associating with PI attacks, threat actors aim for information gathering, fraud, intrusion, malware, content manipulation, and availability attacks choi2022prompt . Carlini, Nicholas, et al carlini2021extracting executed the Data Reconstruction attack on GPT-2, extracting personal identity information, code, and UUIDs. Data Poisoning pertains to the deliberate introduction of corrupted or malicious data into the training dataset to manipulate the model’s behavior. Member Inference Attack indicates that models tend to memorize sensitive information, posing them as potential targets for attacks. A study postulates that an attacker can train a binary classifier using the input-output pairs from LLMs to determine whether a specific individual was part of the model’s training set 7958568 . Diverging from existing works, our research innovatively suggests attacking the capabilities of language models from the perspective of Tokenization.
@@ 3 Methodology of dataset construction
In this section, we describe the process of constructing ADT (Adversarial Dataset for Tokenizer) in detail, which is used to challenge LLMs’ tokenization and thus reveal LLMs’ vulnerability. Our ADT contains two subsets, the manually constructed ADT-Human and automatically generated ADT-Auto.
@@ 3.1 Vocabulary export
In fact, an instance in our dataset comprises one sentence containing the token (word) that could challenge LLMs’ tokenization and one question related to the sentence. To inspect whether LLMs can accurately recognize these challenging tokens within various contexts based on their memories, the tokens should come from the vocabularies of LLMs. Therefore, the first step of our dataset construction is to export the vocabularies of LLMs. Given that Chinese is more complex and challenging than English in terms of tokenization, as detailed in Section 3.2, we pay more attention to the issues in Chinese data.
Specifically, we select five widely-used open-source LLMs which are trained on extensive Chinese corpus, to export their Chinese vocabularies, including Chatglm3-6B zeng2023glm130b , Baichuan2-13B-Chat yang2023baichuan , Yi-34B-Chat ai2024yi , Qwen-7B-Chat bai2023qwen , and Qwen1.5-72B-Chat bai2023qwen . Besides, we export the English vocabularies from three English-based LLMs, i.e., Llama-3-8B-Instruct touvron2023llama , Llama-3-70B-Instruct touvron2023llama , and Mixtral-8x7B-Instruct-v0.1 jiang2024mixtral . Notably, the Chinese vocabulary of Qwen-7B-Chat and Qwen1.5-72B-Chat is the same, so is the English vocabulary of Llama-3-8B-Instruct and Llama-3-70B-Instruct. These LLMs use SentencePiece or tiktoken to train their tokenizers, which are tokenizer tools widely used by various language models.
In the process of exporting the vocabularies, a sequence of straightforward operations is considered. Initially, the tokenizer is decoded to obtain the vocabulary, followed by the removal of leading and trailing spaces from each token. Notably, if SentencePiece is used for tokenization in the training phase, some certain tokens may begin with a special token ‘_’ because SentencePiece treats the input text just as a sequence of Unicode characters, and whitespace is also handled as a normal symbol. To handle the whitespace as a basic token explicitly, SentencePiece first escapes the whitespace with a meta symbol ‘_’ (U+2581)  [^4] https://github.com/google/sentencepiece. Consequently, when exporting the vocabularies for models trained with SentencePiece, we replace ‘_’ with a blank character.
We summarize the exported vocabulary sizes of different LLMs in Table 1.
@@
Language Model Vocabulary Size Vocabulary Size
of Specific Language
Chinese Chatglm3-6B 64,789 30,922
Baichuan2-13B-Chat 125,696 70,394
Yi-34B-Chat 64,000 21,353
Qwen-7B-Chat 151,851 24,953
Qwen1.5-72B-Chat 151,646 24,953
English Llama-3-8B-Instruct 128,257 72,420
Llama-3-70B-Instruct 128,257 72,420
Mixtral-8x7B-Instruct-v0.1 32,000 25,056
@@ 3.2 ADT-Human construction
Based on the exported vocabularies, we manually construct our dataset ADT-Human, to challenge and evaluate the tokenization of different LLMs. The process of manual construction is depicted in Figure 2.
@@
Specifically, we sophisticatedly select eleven tokens from each Chinese vocabulary, and seven tokens for each English vocabulary. Then, for each selected token, we adopt one of the three approaches listed in Table 2, to convert it into a challenging span through inserting a special character span before or (and) after it. These challenging spans would disrupt the conventional tokenization process, thus confusing LLMs. The considerations in the three conversion approaches are introduced as follows.
@@
    Before: A character span $s$ is inserted before the origin token, causing that the concatenation of $s$ and the token’s prefix is just another token (word) existing in the vocabulary, as ‘move’+‘s’$\rightarrow$‘moves’ in Table 2.
@@
    After: A character span $s$ is inserted after the origin token, causing that the concatenation of the token’s suffix and $s$ is just another token existing in the vocabulary, as ‘n’+‘othing’$\rightarrow$‘nothing’ in Table 2.
@@
    Before & After: Character spans $s_{1}$ and $s_{2}$ are inserted before and after the origin token, respectively, causing that the concatenation of $s_{1}$ and the token’s prefix is just another token existing in the vocabulary, meanwhile so is the concatenation of the token’s suffix and $s_{2}$, as ‘six’+‘th’$\rightarrow$‘sixth’ and ‘is’+‘land’$\rightarrow$‘island’ in Table 2.
Next, for each challenging span $s_{c}$, we manually compose a corresponding instance as the data in ADT-Human. One instance consists of a sentence in which $s_{c}$ presents, and a corresponding question of which the answer comes from $s_{c}$. In Appendix A, we list all instances (along with their correct tokenization) in ADT-Human for the four Chinese vocabularies and two English vocabularies we export. As we can see, it is challenging for the LLMs to understand the instances due to the presence of challenging spans.
Table 2: Three approaches of generating challenging spans.
Approach Origin token After insertion Model tokenization Human tokenization
(Challenging span)
Before Schema BC ABC [A, BC] [AB, C]
Example stable movestable [move, stable] [moves, table]
After Schema AB ABC [AB, C] [A, BC]
Example driven drivenothing [driven, othing] [drive, nothing]
Before & After Schema BC ABCD [A, BC, D] [AB, CD]
Example this sixthisland [six, this, land] [sixth, island]
It’s worth noting that the tokenization difficulty of English is less than that of Chinese, since spaces are regularly used as delimiters to separate each word from others in English. Moreover, affixation is common in English word structure, enabling tokenizers to divide a single word into several sections, which can help avoid incorrect tokenization to some extent. Thus, during the manual construction of English instances, we deliberately exclude the spaces between some tokens to provoke challenges to tokenization process.

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable capabilities in language understanding and generation. Nonetheless, it was also witnessed that LLMs tend to produce inaccurate responses to specific queries. This deficiency can be traced to the tokenization step LLMs must undergo, which is an inevitable limitation inherent to all LLMs. In fact, incorrect tokenization is the critical point that hinders LLMs in understanding the input precisely, thus leading to unsatisfactory output. This defect is more obvious in Chinese scenarios. To demonstrate this flaw of LLMs, we construct an adversarial dataset, named as ADT (Adversarial Dataset for Tokenizer), which draws upon the vocabularies of various open-source LLMs to challenge LLMs’ tokenization. ADT consists of two subsets: the manually constructed ADT-Human and the automatically generated ADT-Auto. Our empirical results reveal that our ADT is highly effective on challenging the tokenization of leading LLMs, including GPT-4o, Llama-3, Deepseek-R1 and so on, thus degrading these LLMs’ capabilities. Moreover, our method of automatic data generation has been proven efficient and robust, which can be applied to any open-source LLMs. In this paper, we substantially investigate LLMs’ vulnerability in terms of challenging their token segmentation, which will shed light on the subsequent research of improving LLMs’ capabilities through optimizing their tokenization process and algorithms.
Tokenization Matters! Degrading Large Language Models through Challenging Their Tokenization
Dixuan Wang1, Yanda Li1, Junyuan Jiang2, Zepeng Ding1, Ziqin Luo1, Guochao Jiang1, Jiaqing Liang1thanks: Corresponding Authors, Deqing Yang [^1] footnotemark: 1 School of Data Science, Fudan University, Shanghai, China 1 School of Management, Fudan University, Shanghai, China 2 {dxwang23,ydli22,jiangjy21,zqluo22,gcjiang22}@m.fudan.edu.cn, {dingzepeng,liangjiaqing,yangdeqing}@fudan.edu.cn
@@ 1 Introduction
In the last two years, Large Language Models (LLMs) have demonstrated remarkable capabilities in many tasks of artificial intelligence (AI), including natural language generation Hoffmann et al. (2022); Nijkamp et al. (2023); Zeng et al. (2023), knowledge utilization Chowdhery et al. (2023); Izacard et al. (2023), and complex reasoning Wei et al. (2022); Zhou et al. (2023); Kojima et al. (2022); Taylor et al. (2022). Given these capabilities, LLMs have been effectively employed by various application domains, such as healthcare Tang et al. (2023); Li et al. (2023); Jeblick et al. (2022), education Susnjak (2022); Bordt and von Luxburg (2023); Malinka et al. (2023), law Yu et al. (2022); Nay (2022) and so on.
@@
Figure 1: Two instances of LLM generating incorrect response due to incorrect tokenization. Case 1 is a Chinese input instance, of which the English translation is noted below according to its correct tokenization. In Case 2, a space is omitted between ‘sing’ and ‘lemon’, causing the LLM’s incorrect tokenization, which is detailed in Section 3.2.
Nonetheless, LLMs’ disadvantages have also been witnessed, including hallucination OpenAI (2023); Bang et al. (2023); Lin et al. (2022), knowledge recency Dai et al. (2022); Kernbach (2022), and so on. Particularly, we observed that for some specific queries, LLMs often produce unsatisfactory responses with words that are nonsensical, as illustrated by the two instances in Figure 1. Through checking the LLM’s tokenization results for input sentences, we found that it is misaligned with human’s correct comprehension for the sentences. Notably, our empirical studies found that this flaw not only exists in some specific LLMs, but also is a universal issue across many mainstream LLMs. We have evaluated several prominent open-source and closed-source LLMs, including Chatglm3 Zeng et al. (2023), Qwen2.5-max Team (2024), Deepseek-R1 DeepSeek-AI (2025), and GPT-4o OpenAI (2024). Our experiment results reveal that regardless of LLMs’ scales or their claimed capabilities, they inevitably generate incorrect or entirely nonsensical responses for some specific inputs when their tokenization results for the input sentences are obviously wrong. Consequently, we believe that LLMs’ tokenization errors prevent them from accurately understanding the input text, leading to their incorrect responses.
@@
As we know, LLMs’ tokenization flaws stem from the algorithms of their tokenizers, most of which are based on subword-level vocabularies. The popular tokenization algorithms include Byte-Pair Encoding (BPE) Sennrich et al. (2016), WordPiece Schuster and Nakajima (2012), and Unigram Kudo (2018). However, no vocabulary can perfectly cover all possible ways of various expressions in the inputs. The algorithms may potentially generate unsatisfactory results which are not fully aligned with the true intention of users’ input. Unfortunately, in cases of tokenization errors, all subsequent optimization operations for LLMs cannot completely solve this underlying problems caused by their tokenization algorithms.
@@
In the domain of natural language processing (NLP), the existing studies related to tokenization primarily focus on refining or optimizing various tokenization algorithms. Meanwhile, the discussions on LLMs’ vulnerability including attack or challenge techniques, have been more concerned with the security of LLMs. For LLMs, in terms of the challenges posed by tokenization deficiencies, Sander and Max Land and Bartolo (2024) have discussed this issue from the perspective of under-trained tokens in LLMs. It is worth noting that, to the best of our knowledge, there has been no research specifically focusing on the unsatisfactory token segmentation of LLMs, particularly in Chinese scenarios, which is indeed a critical concern causing LLMs’ vulnerability.
@@
In this paper, we focus on the critical flaw in LLMs’ tokenization process, and try to reveal the relationship between LLMs’ unsatisfactory tokenization and their inaccurate responses for some specific queries. To this end, for the first time, we construct a dataset, namely ADT (Adversarial Dataset for Tokenizer), to challenge the tokenization of various LLMs. ADT dataset consists of two subsets: the manually constructed ADT-Human and the automatically generated ADT-Auto. At first, we export the vocabularies from multiple mainstream LLMs, based on which ADT-Human is constructed. Our experiment results demonstrate that ADT-Human can effectively challenge LLMs’ tokenization, leading to their completely incorrect responses. Furthermore, we also develop a framework for automatically generating adversarial data to construct dataset more efficiently. Initially, we export LLM’s vocabulary and identify the trap words that can influence the model’s performance. By inputting these trap words into GPT-4 OpenAI (2023) with prompt, we leverage its capability to get available instances, which are then inspected manually to ensure quality. With minimal human effort, we successfully construct ADT-Auto with 231 instances, validating the effectiveness of our framework and highlighting the inevitable flaws in LLMs’ tokenization once again.
@@
    We investigate LLMs’ vulnerability for some special inputs in terms of challenging their token segmentation, which provides a new perspective of studying LLMs’ disadvantages.
@@
    Our experiment results obviously reveal the relationship between LLMs’ unsatisfactory tokenization and inaccurate responses, which can shed light on subsequent work of improving LLMs through optimizing tokenization.
@@ 2 Related Work
Tokenization is a basic but crucial step in NLP. Currently, the mainstream tokenization approachs are based on subwords, including Byte Pair Encoding (BPE), WordPiece and Unigram. BPE Sennrich et al. (2016) forms the vocabulary starting from character-level tokens, merging token pairs and intercalating them into vocabulary. The merging rule is to select adjacent pairs with the highest word frequency. LLMs including GPT-3 Brown et al. (2020), RoBERTa Liu et al. (2019), and Llama2 Touvron et al. (2023) are based on BPE. WordPiece Schuster and Nakajima (2012) is similar to BPE, but it differs in the strategy of merging pairs with reference to mutual information rather than frequency. LLMs built upon WordPiece include BERT Devlin et al. (2019), DistilBERT Sanh et al. (2019), and Electra Clark et al. (2020). Different from these two algorithms, Unigram Kudo (2018) starts with a large vocabulary and gradually trims it down to a smaller one. It measures the importance of subwords by calculating loss associated with the removal of each subword, ultimately retaining those exhibit high importance. LLMs utilizing Unigram include AlBERT Lan et al. (2020) and mBART Liu et al. (2020). Besides, many tokenization tools integrating these algorithms have been developed, such as Google’s SentencePiece Kudo and Richardson (2018) and OpenAI’s tiktoken [^1] https://github.com/openai/tiktoken, which simplify the process of LLMs tokenization greatly. In recent years, it has been investigated how the input segmentation of pre-trained language models (PLMs) affects the interpretations of derivationally complex English words Hofmann et al. (2021). Some scholars have proposed FLOTA Hofmann et al. (2022), a simple yet effective method to improve PLMs’ tokenization of English words. However, there has been no work concerning the shared risks of LLMs in terms of inaccurate tokenization result. We focus on this issue, discussing the underlying risks of LLMs’ tokenization.
@@
With the growing prominence of LLMs, the security and vulnerability of these models have attracted significant attention, and even advanced LLMs like GPT-4 are no exception. A surge of research in this field is underway, with researchers launching attacks on LLMs from various aspects Esmradi et al. (2023); Chowdhury et al. (2024) including but not limited to, Prompt Injection Choi et al. (2022), Model Theft Krishna et al. (2020), Data Reconstruction Carlini et al. (2021), Data PoisoningWallace et al. (2021); Xu et al. (2023), and Member Inference Attack Liu et al. (2023). For instance, Prompt Injection Attack refers to a scenario where an attacker crafts malicious prompts to deceive language models into generating outputs inconsistent with their training data and anticipated functionality. Threat actors aim at information gathering, fraud, intrusion, content manipulation, and availability attacks Choi et al. (2022). Carlini, Nicholas, et al Carlini et al. (2021) executed the Data Reconstruction attack on GPT-2, extracting personal identity information, code, and UUIDs. Data Poisoning pertains to the deliberate introduction of corrupted or malicious data into the training dataset to manipulate the model’s behavior. Diverging from existing works, our research innovatively suggests attacking the capabilities of LLMs from the perspective of tokenization.
@@ 3 Methodology of Dataset Construction
In this section, we describe the process of constructing ADT (Adversarial Dataset for Tokenizer) in detail, which is used to challenge LLMs’ tokenization and thus reveal LLMs’ vulnerability. ADT contains two subsets, manually constructed ADT-Human and automatically generated ADT-Auto.
@@ 3.1 Vocabulary Export
In fact, an instance in our dataset comprises one sentence containing the token (word) that could challenge LLMs’ tokenization and one question related to the sentence. To inspect whether LLMs can accurately recognize these challenging tokens within various contexts based on their memories, the tokens should come from vocabularies of LLMs. Therefore, the first step of dataset construction is to export vocabularies of LLMs. Given that Chinese is more complex and challenging than English in terms of tokenization, as detailed in Section 3.2, we pay more attention to the issue in Chinese data.
Specifically, we select five widely-used open-source LLMs which are trained on extensive Chinese corpus, to export their Chinese vocabularies, including Chatglm3-6B Zeng et al. (2023), Baichuan2-13B-Chat Yang et al. (2023), Yi-34B-Chat Young et al. (2024), Qwen-7B-Chat Bai et al. (2023), and Qwen1.5-72B-Chat Bai et al. (2023). Besides, we export the English vocabularies from three English-based LLMs, i.e., Llama-3-8B-Instruct Touvron et al. (2023), Llama-3-70B-Instruct Touvron et al. (2023), and Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024). Notably, the Chinese vocabulary of Qwen-7B-Chat and Qwen1.5-72B-Chat is the same, so is the English vocabulary of Llama-3-8B-Instruct and Llama-3-70B-Instruct.
Language Model Vocabulary Size Vocabulary Size of Specific Language Chinese Chatglm3-6B 64,789 30,922 Baichuan2-13B-Chat 125,696 70,394 Yi-34B-Chat 64,000 21,353 Qwen-7B-Chat 151,851 24,953 Qwen1.5-72B-Chat 151,646 24,953 English Llama-3-8B-Instruct 128,257 72,420 Llama-3-70B-Instruct 128,257 72,420 Mixtral-8x7B-Instruct-v0.1 32,000 25,056
@@
In the process of exporting vocabularies, a sequence of straightforward operations is considered. Initially, the tokenizer is decoded to obtain vocabulary, followed by the removal of leading and trailing spaces from each token. Notably, if SentencePiece is used for tokenization in training phase, some certain tokens may begin with a special token ‘_’ because SentencePiece treats the input text just as a sequence of Unicode characters, and whitespace is also handled as a normal symbol. To handle the whitespace as a basic token explicitly, SentencePiece first escapes the whitespace with a meta symbol ‘_’ (U+2581) Kudo and Richardson (2018). Consequently, when exporting vocabularies for models trained with SentencePiece, we replace ‘_’ with a blank character. We summarize the exported vocabulary sizes of different LLMs in Table 1.
@@ 3.2 ADT-Human Construction
Based on exported vocabularies, we manually construct dataset ADT-Human, to challenge and evaluate the tokenization of different LLMs. The process of manual construction is depicted in Figure 2.
@@
Approach Origin token After insertion Model tokenization Human tokenization (Challenging span) Before Schema BC ABC [A, BC] [AB, C] Example stable movestable [move, stable] [moves, table] After Schema AB ABC [AB, C] [A, BC] Example driven drivenothing [driven, othing] [drive, nothing] Before & After Schema BC ABCD [A, BC, D] [AB, CD] Example this sixthisland [six, this, land] [sixth, island]
Table 2: Three approaches of generating challenging spans.
Our purpose is to confirm the existence of challenges in tokenization, so for each LLM, a certain amount of data is constructed, which does not need to be large. Specifically, we select eleven tokens from each Chinese vocabulary, and seven tokens for each English vocabulary. The main criterion for selecting tokens is that they should be easy to make sentences with. According to the experimental results presented in Section 4.2, these data can effectively challenge the performance of LLMs, proving this issue deserves more attention. As for efficiently generating data in bulk, we design the framework for automatic generation in Section 3.3.
Then, for each selected token, we adopt one of the three approaches listed in Table 2, to convert it into a challenging span through inserting a special character span before or (and) after it. These challenging spans would disrupt the conventional tokenization process, thus confusing LLMs. The schemas and examples of three approaches are also shown in the second and third parts of Figure 2. The considerations in the three conversion approaches are introduced as follows.
@@
    Before: A character span $s$ is inserted before origin token, causing the concatenation of $s$ and the token’s prefix is just another token existing in vocabulary, as ‘move’+‘s’$\rightarrow$‘moves’ in Table 2.
@@
    After: A character span $s$ is inserted after origin token, causing the concatenation of the token’s suffix and $s$ is just another token existing in vocabulary, as ‘n’+‘othing’$\rightarrow$‘nothing’ in Table 2.
@@
    Before & After: Character spans $s_{1}$ and $s_{2}$ are inserted before and after origin token, respectively, causing the concatenation of $s_{1}$ and the token’s prefix is just another token existing in vocabulary, meanwhile so is the concatenation of the token’s suffix and $s_{2}$, as ‘six’+‘th’$\rightarrow$‘sixth’ and ‘is’+‘land’$\rightarrow$‘island’ in Table 2.
Next, for each challenging span $s_{c}$, we manually compose a corresponding instance as the data in ADT-Human. One instance consists of a sentence in which $s_{c}$ presents, and a corresponding question of which the answer comes from $s_{c}$. In Appendix A, we list all instances (along with their correct tokenizations) of ADT-Human. As we can see, it is challenging for LLMs to understand the instances due to the presence of challenging spans.
It’s worth noting that the tokenization difficulty of English is less than that of Chinese, since spaces are regularly used as delimiters to separate each word from others in English. Moreover, affixation is common in English word structure, enabling tokenizers to divide a single word into several sections, which can help avoid incorrect tokenization to some extent. Thus, during the manual construction of English instances, we deliberately exclude the spaces between some tokens to provoke challenges to tokenization process. This decision stems from the recognition that powerful models should possess robust abilities across various scenarios which can occur in real-world applications, including handling cases where spaces might be omitted in English text input Hofmann et al. (2022).